How many liters each of a 35% acid solution and a 85% acid solution must be used to produce 50 liters of a 55% acid solution? (Round to two decimal places if necessary.)
Answer by TimothyLamb(4379)

x = liters of 35% acid
y = liters of 85% acid
---
x + y = 50
35x + 85y = 55*50
---
put the system of linear equations into standard form
---
x + y = 50
35x + 85y = 2750
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x = liters of 35% acid = 30
y = liters of 85% acid = 20
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---